OpenStudy (anonymous):

The radius R of a sphere is 7.3m Calcualte the sphere's surface area A. use the value 3.14 for pie, and round your answer to the nearest tenth.

OpenStudy (anonymous):

surface area = 4pr2 r=7.3m surface area=669.32m^2

OpenStudy (anonymous):

\[surface area =4 \pi r^{2}\] plug in the values given\[area =4 \times 3.14 \times 7.3^{2}\] \[area=669.3224m^{2}\] round to the nearest tenth \[area =669.3 m^{2}\]

OpenStudy (anonymous):

what would the answer be if rounded to the nearest tenth

OpenStudy (anonymous):

tenth is the the first place after the decimal point

OpenStudy (anonymous):

ohh! ok thankyou the answer is right

OpenStudy (anonymous):

wanna help me with another one

OpenStudy (anonymous):

depends on the question post it

OpenStudy (anonymous):

aight i posted it there:D